CAVITY RINGDOWN SPECTROSCOPY OF THE NU1 BAND OF HNO3: DERIVING ACCURATE INTE-GRATED INTENSITIES FROM BROAD BAND MEASUREMENTS

\underline{ANDREW K. MOLLNER}, GAUTHAM NAIR, MITCHIO OKUMURA, Department of Chemistry, Caltech, Pasadena, CA.

Cavity ringdown spectroscopy (CRDS) is a method which utilizes a highly reflective optical cavity to obtain high spectroscopic sensitivity. Given its sensitivity, it is desirable to use CRDS to make quantitative measurements of trace species. We have conducted a theoretical and experimental analysis of the ν1 band of HNO3 to evaluate the accuracy of our infrared-CRDS apparatus as a quantitative spectroscopic tool.